OpenStudy (anonymous):

Sulfur vapor is analyzed by photoelectron spectroscopy (PES). Measurements determine that photoelectrons associated with the 1st ionization energy of sulfur move with de Broglie wavelength λ=3.091 A˚. What is the maximum wavelength (in meters) of radiation capable of ionizing sulfur and producing this effect?

OpenStudy (anonymous):

did you get an answer?

OpenStudy (anonymous):

Can't get it either

OpenStudy (anonymous):

i got the ans... they hv given the debrogli wavelength so....velocity is 2.353*10^6m/s using lambda=h/mv and E=25.21*10^-19J using E=1/2mv2 thn h*nu=K.E+B.E here B.E=I.E of s in J wch is 10.36eV*1.602*10^-19 =16.5967*10^-19 nw using eq h*nu=K.E+B.E i gt lambda.. =47.4*10^-9